# Asymptotic Normality of Subgraph Counts in Sparse Inhomogeneous Random Graphs

### v1

@@ 2.4.3. Dense Graphon Model
Proof of Proposition 6.1.
Recall the definition of $T(H,G_{n})$ from (5.1). Since $W$ is not $H$-regular, by Lemma 5.1 $T(H,G_{n})$ is an $U$-statistic of order $|V(H)|$ which is non-degenerate. Hence, by [37, Section 1.3 Theorem 3], $\mathrm{Var}[T(H,\,G_{n})]\lesssim_{H,W}\frac{1}{n}$, and recalling (5.2),
Also, recalling (3.2),
$\mathrm{Var}[\Delta(H,G_{n})]\asymp_{H,W}\max_{F\subseteq H:|V(F)|\geq 1}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}.$ (6.1)
We claim that when $n\rho_{n}^{m_{1}(H)}\ll 1,$ the maximum in (6.1) is obtained for some subgraph $F\subseteq H$ with $|V(F)|\geq 2.$ Indeed, when $n\rho_{n}^{m_{1}(H)}\ll 1,$ there exists subgraph $F\subseteq H$ with $|V(F)|\geq 2$ such that $n^{|V(F)|-1}\rho_{n}^{|E(F)|}\ll 1$ or equivalently, $n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}\ll n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}.$ Thus,
$\frac{\mathrm{Var}[\Delta_{2}(H,\,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\lesssim_{H,\,W}\min_{F\subseteq H,\,|V(F)|\geq 2}n^{|V(F)|-1}\rho_{n}^{|E(F)|}\to 0$
since $n\rho_{n}^{m_{1}(H)}\ll 1.$ This proves Proposition 6.1 (1).
Next, for the second part, from [37, Section 1.3, Theorem 3] we know that
$\lim_{n\to\infty}n\mathrm{Var}[T(H,\,G_{n})]=|V(H)|^{2}\xi_{1},$
$\mathrm{Var}[\Delta_{2}(H,\,G_{n})]=\frac{|V(H)|^{2}}{n}\binom{n}{|V(H)|}^{2}\rho_{n}^{2|E(H)|}\xi_{1}+o(n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}).$ (6.2)
Now, continuing from (3.4), we have
$\mathrm{Var}[\Delta(H,\,G_{n})]$ $=\sum_{F\subseteq H:|V(F)|=1}\sum_{\stackrel{{\scriptstyle H_{s},\,H_{t}\in\mathcal{H}_{n}}}{{H_{s}\cap H_{t}\simeq F}}}\mathrm{Cov}[I_{s},\,I_{t}]+\sum_{F\subseteq H:|V(F)|\geq 2}\sum_{\stackrel{{\scriptstyle H_{s},\,H_{t}\in\mathcal{H}_{n}}}{{H_{s}\cap H_{t}\simeq F}}}\mathrm{Cov}[I_{s},\,I_{t}],$ (6.3)
where $F$ is the subgraph of $H$ isomorphic to $H_{s}\cap H_{t}=(V(H_{s})\cap V(H_{t}),\,E(H_{s})\cap E(H_{t}))$. Note that for a fixed subgraph $F\subseteq H,$ the number of pairs $(H_{s},\,H_{t})$ having $H_{s}\cap H_{t}$ isomorphic to $F$ is $O_{H}(n^{2|V(H)|-|V(F)|}).$ Hence, recalling (3.1), the second term in (6.3) can be bounded as
$\sum_{F\subseteq H:|V(F)|\geq 2}\sum_{\stackrel{{\scriptstyle H_{s},\,H_{t}\in\mathcal{H}_{n}}}{{H_{s}\cap H_{t}\simeq F}}}\mathrm{Cov}[I_{s},\,I_{t}]$ $\lesssim_{H,\,W}\sum_{F\subseteq H:|V(F)|\geq 2}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}$
$=o(n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}),$
since $n\rho_{n}^{m_{1}(H)}\gg 1$ is equivalent to the condition that for all subgraph $F\subseteq H$ with $|V(F)|\geq 2,$ we have $n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}\ll n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}.$ To compute the first term in (6.3), first choose the vertex sets of the two copies of $H$ sharing a single vertex. This is equivalent to finding the number of ordered pairs of subsets $(A_{1},\,A_{2})$ of $[n]$ such that $|A_{1}\cap A_{2}|=1.$ Observe that the number of such ordered pairs is $|V(H)|\binom{n}{|V(H)|}\binom{n-|V(H)|}{|V(H)|-1}$. Once we have fixed the vertex sets of the two copies, we can trace the steps of the proof of Lemma 5.1 to obtain
$\mathrm{Var}[\Delta(H,G_{n})]=|V(H)|\binom{n}{|V(H)|}\binom{n-|V(H)|}{|V(H)|-1}\rho_{n}^{2|E(H)|}\xi_{1}+o(n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}).$ (6.4)
from (6.3). Taking the ratio of (6.2) and (6.4) it follows that $\frac{\mathrm{Var}[\Delta_{2}(H,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\rightarrow 1$. Hence, recalling (2.3) the result in Proposition 6.1 (2) follows. ∎
Using the above result, we now complete the proof of Theorem 2.11. To begin with suppose, $n\rho_{n}^{m_{1}(H)}\ll 1$. The first assertion in (2.8) follows from Proposition 6.1 (1). This means, using (2.3), $\frac{\mathrm{Var}[\Delta_{1}(H,\,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}=1-\frac{\mathrm{Var}[\Delta_{2}(H,\,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\to 1$. Hence, from (4.11), $\frac{\mathrm{Var}[\Delta_{1}(H,\,G_{n})|\mathcal{F}(\bm{U}_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\rightarrow 1$ almost surely. Hence, applying Proposition 2.15 and Slutsky’s theorem it follows that
@@
Next, suppose $n\rho_{n}^{m_{1}(H)}\gg 1$. The first assertion in (2.9) follows from Proposition 6.1 (2). Also, since $W$ is not $H$-regular, by Lemma 5.1 $\Delta_{2}(H,\,G_{n})$ is a scaled $U$-statistic of order $|V(H)|$ which is non-degenerate. Hence, by [37, Chapter 3, Theorem 1],
@@
Next, consider the case $n\rho_{n}^{m_{1}(H)}\to c,$ for some $c\in(0,\,\infty).$ We have, from (6.2),
@@
$\mathrm{Var}[\Delta_{1}(H,\,G_{n})]$ $=\sum\limits_{\begin{subarray}{c}F\subseteq H\\ |E(F)|\geq 1\end{subarray}}\left\{\rho_{n}^{2|E(H)|-|E(F)|}\mathbb{E}[Z_{n}(F)]+o(n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|})\right\}$ (6.6)
@@
To find the dominating term in above sum, define the following set:
@@
Then applying Slutsky’s theorem with (6.10) proves the result in (2.10). $\Box$
@@
recalling (1.2). Here, $|\mathscr{G}_{H}|$ is the number of copies of $H$ in the complete graph $K_{|V(H)|}$ on $|V(H)|$ vertices, and $\xi_{1}$ is defined in Appendix A. This shows that under $n\rho_{n}^{m_{1}(H)}\to c\in(0,\,\infty),$
@@
where $\kappa\in(0,\,1)$ is a constant depending on c,Hc,\,H and $W.$
@@ 7. Conclusions
In this paper, we show that $X(H,G_{n})$, with $G_{n}\sim G(n,\rho_{n},W)$, is asymptotically normal whenever $n\rho_{n}^{m(H)}\gg 1$, that is, when the sparsity level is above the containment threshold for $H$. We do not consider the critical regime $n\rho_{n}^{m(H)}\asymp 1$. For the classical Erdős–Rényi model, it is known that at the threshold, $X(H,G_{n})$ converges to a Poisson distribution when $H$ is strictly balanced [10]. If $H$ is unbalanced, the problem reduces to a balanced subgraph via an appropriate normalization [52, Section 4]. When $H$ is balanced but not strictly balanced, the limiting distribution is more intricate and depends on the structure and multiplicity of subgraphs maximizing $m(H)$, with no universal closed form [11, 34].
Recently, there has been growing interest in deriving analogous results for inhomogeneous random graph models. In this direction, Coulson et al. [21] derived Poisson approximation results for $X(H,G_{n})$, when $G_{n}$ is sampled from a stochastic block model and $H$ is strictly balanced (see also [22]). Recently, Liu and Privault [39] obtained an analogous Poisson approximation result in the random connection model. Extending these results to the sparse graphon model, as well as addressing the case where $H$ is not strictly balanced, are interesting directions for future research.
@@ Appendix A Proofs of Technical Lemmas
where the first $|V(F)|$ vertices are common. The edge sets of $H_{s},\,H_{t}$ are determined by maps $\phi_{s}:V(H)\to V(H_{s})$ and $\phi_{t}:V(H)\to V(H_{t})$, such that $(a,\,b)\in E(H)$ if and only if $(\phi_{s}(a),\,\phi_{s}(b))\in E(H_{s})$ and $(\phi_{t}(a),\,\phi_{t}(b))\in E(H_{t}).$ In particular, we consider the following maps
@@
is the $|V(F)|$-point conditional homomorphism density of $H$ in $W$ (see [5, Definition 2.1]). Then, by the Cauchy-Schwarz inequality,
[…]
@@
Furthermore, note that to create pairs $H_{s},\,H_{t}$ as above, we may first choose the common $|V(F)|$ vertices in at most $\binom{n}{|V(F)|}$ ways. The rest of the vertices of $H_{s}$ and $H_{t}$ can be chosen in at most
@@
ways. Hence, the number of such pairs is $O_{H}(n^{2|V(H)|-|V(F)|})$. This completes the proof of the lemma. $\Box$
@@
$\xi_{1}$ $=\mathrm{Cov}(h(U_{1},\,U_{2},\,\ldots,\,U_{|V(H)|}),\,h(U_{|V(H)|},\,U_{|V(H)|+1},\,\ldots,\,U_{2|V(H)|-1}))$
$=\mathbb{E}(h(U_{1},\,U_{2},\,\ldots,\,U_{|V(H)|})h(U_{|V(H)|},\,U_{|V(H)|+1},\,\ldots,\,U_{2|V(H)|-1}))$
@@
$=\sum\limits_{\begin{subarray}{c}H^{\prime}\in\mathscr{G}_{H}(\{1,\,2,\,\ldots,\,|V(H)|\})\\ H^{\prime\prime}\in\mathscr{G}_{H}(\{|V(H)|,\,|V(H)|+1,\,\ldots,\,2|V(H)|-1\})\end{subarray}}\left[\mathbb{E}\left(\prod\limits_{\begin{subarray}{c}(i,\,j)\in E(H^{\prime})\\ (k,\,l)\in E(H^{\prime\prime})\end{subarray}}W(U_{i},\,U_{j})W(U_{k},\,U_{l})\right)-t(H,\,W)^{2}\right].$
To further simplify the expression we recall [5, Definition 2.7]: For $a,b\in\{1,2,\ldots,|V(H)|\}$ the $(a,b)$-vertex join of two copies of $H$ is the graph obtained by identifying the $a$-th vertex of first copy of $H$ with the $b$-th vertex of the second copy of $H$. The resulting graph is denoted by $H\displaystyle{\bigoplus_{a,b}}H$. Now, applying [5, Lemma 5.2] gives,
$\xi_{1}=\frac{|\mathscr{G}_{H}|^{2}}{|V(H)|^{2}}\left[\sum_{1\leq a,\,b\leq|V(H)|}t\left(H\bigoplus_{a,\,b}H,\,W\right)-|V(H)|^{2}(t(H,\,W))^{2}\right].$
@@
Finally, invoking [5, Lemma 5.4] we conclude that $\xi_{1}=0$ if and only if $W$ is $H$-regular. Thus, when $W$ is $H$-regular, $T(H,G_{n})$ is a first order degenerate $U$-statistics. Hence by [37, Section 1.3, Theorem 3], $\mathrm{Var}[T(H,\,G_{n})]\lesssim_{H,\,W}\frac{1}{n^{2}}$. This completes the proof of Lemma 5.1. $\Box$
A.4. Joint Convergence from Marginal Convergence

### v2

@@ 2.4.3. Dense Graphon Model
The proof of Proposition 6.1 is given in Appendix A.4. Using the above result, we now complete the proof of Theorem 2.11. To begin with, suppose $n\rho_{n}^{m_{1}(H)}\ll 1$. The first assertion in (2.8) follows from Proposition 6.1 (1). This means, using (2.3), $\frac{\mathrm{Var}[\Delta_{1}(H,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}=1-\frac{\mathrm{Var}[\Delta_{2}(H,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\to 1$. Hence, from (4.11), $\frac{\mathrm{Var}[\Delta_{1}(H,G_{n})|\mathcal{F}(\bm{U}_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\rightarrow 1$ almost surely. Hence, applying Proposition 2.15 and Slutsky’s theorem it follows that
@@
Next, suppose $n\rho_{n}^{m_{1}(H)}\gg 1$. The first assertion in (2.9) follows from Proposition 6.1 (2). Also, since $W$ is not $H$-regular, by Lemma 5.1 $\Delta_{2}(H,G_{n})$ is a scaled $U$-statistic of order $|V(H)|$ which is non-degenerate. Hence, by Lee (2019, Chapter 3, Theorem 1),
@@
Next, consider the case $n\rho_{n}^{m_{1}(H)}\to c,$ for some $c\in(0,\infty).$ We have, from (A.2),
@@
$\hskip-3.61371pt\mathrm{Var}[\Delta_{1}(H,G_{n})]$ $=\sum\limits_{\begin{subarray}{c}F\subseteq H\\ |E(F)|\geq 1\end{subarray}}\left\{\rho_{n}^{2|E(H)|-|E(F)|}\mathbb{E}[Z_{n}(F)]+o(n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|})\right\}$ (6.2)
@@
To find the dominating term in the above sum, define the following set:
@@
Then applying Slutsky’s theorem with (6.6) proves the result in (2.10).
$\Box$
@@
recalling (1.2). Here, $|\mathscr{G}_{H}|$ is the number of copies of $H$ in the complete graph $K_{|V(H)|}$ on $|V(H)|$ vertices, and $\xi_{1}$ is defined in Appendix A.3. This shows that under $n\rho_{n}^{m_{1}(H)}\to c\in(0,\infty),$
@@
where $\kappa\in(0,1)$ is a constant depending on $c,H$ and $W.$
@@ 7. Conclusions
In this paper, we show that $X_{n}(H,G_{n})$, with $G_{n}\sim G(n,\rho_{n},W)$, is asymptotically normal whenever $n\rho_{n}^{m(H)}\gg 1$, that is, when the sparsity level is above the containment threshold for $H$. Another important regime which we do not consider in the paper is when $n\rho_{n}^{m(H)}\asymp 1$, that is, $\rho_{n}$ is at the containment threshold. For the classical Erdős–Rényi model, it is known that at the threshold, $X_{n}(H,G_{n})$ converges to a Poisson distribution when $H$ is strictly balanced (Bollobás, 1981). If $H$ is unbalanced, the problem reduces to a balanced subgraph via an appropriate normalization (Rucinski, 1990, Section 4). When $H$ is balanced but not strictly balanced, the limiting distribution is more intricate and depends on the structure and multiplicity of subgraphs maximizing $m(H)$, with no universal closed form (Bollobás and Wierman, 1989, Janson et al., 2011).
Recently, there has been growing interest in deriving analogous results for inhomogeneous random graph models. In this direction, Coulson et al. (2016) derived Poisson approximation results for $X_{n}(H,G_{n})$, when $G_{n}$ is sampled from a stochastic block model and $H$ is strictly balanced (see also Coulson et al. (2018)). Recently, Liu and Privault (2025) obtained an analogous Poisson approximation result in the random connection model. Extending these results to the sparse graphon model, as well as addressing the case where $H$ is not strictly balanced, when $\rho_{n}$ is at the containment threshold, are interesting directions for future research.
@@ Appendix A Proofs of Technical Lemmas
where the first $|V(F)|$ vertices are common. The edge sets of $H_{s},\,H_{t}$ are determined by maps $\phi_{s}:V(H)\to V(H_{s})$ and $\phi_{t}:V(H)\to V(H_{t})$, such that $(a,\,b)\in E(H)$ if and only if $(\phi_{s}(a),\,\phi_{s}(b))\in E(H_{s})$ and $(\phi_{t}(a),\,\phi_{t}(b))\in E(H_{t})$, respectively. In particular, we consider the following maps
@@
is the $|V(F)|$-point conditional homomorphism density of $H$ in $W$ (see Bhattacharya et al. (2023, Definition 2.1)). Then, by the Cauchy-Schwarz inequality and using $|W|\leq 1$,
$t(H_{s}\cup H_{t},\,W)$ $\geq\left(\int_{[0,\,1]^{|V(F)|}}t_{V(F)}(x_{1},\,x_{2},\,\ldots,\,x_{|V(F)|};\,H,\,W)\prod_{a\in V(F)}\text{d}x_{a}\right)^{2}$
@@
Furthermore, note that to create pairs $H_{s},\,H_{t}$ as above, we may first choose the common $|V(F)|$ vertices in $\binom{n}{|V(F)|}$ ways. The rest of the vertices of $H_{s}$ and $H_{t}$ can be chosen in
@@
ways. Hence, the number of such pairs is $\Theta_{H}(n^{2|V(H)|-|V(F)|})$. This completes the proof of the lemma. $\Box$
@@
$\xi_{1}$ $=\mathrm{Cov}[h(U_{1},\,U_{2},\,\ldots,\,U_{|V(H)|}),\,h(U_{|V(H)|},\,U_{|V(H)|+1},\,\ldots,\,U_{2|V(H)|-1})]$
$=\mathbb{E}[h(U_{1},\,U_{2},\,\ldots,\,U_{|V(H)|})h(U_{|V(H)|},\,U_{|V(H)|+1},\,\ldots,\,U_{2|V(H)|-1})]$
@@
$=\sum\limits_{\begin{subarray}{c}H^{\prime}\in\mathscr{G}_{H}(\{1,\,2,\,\ldots,\,|V(H)|\})\\ H^{\prime\prime}\in\mathscr{G}_{H}(\{|V(H)|,\,|V(H)|+1,\,\ldots,\,2|V(H)|-1\})\end{subarray}}\left\{\mathbb{E}\left[\prod\limits_{\begin{subarray}{c}(i,\,j)\in E(H^{\prime})\\ (k,\,l)\in E(H^{\prime\prime})\end{subarray}}W(U_{i},\,U_{j})W(U_{k},\,U_{l})\right]-t(H,\,W)^{2}\right\}.$
To further simplify the expression recall Definition 3.3. Now, applying (Bhattacharya et al., 2023, Lemma 5.2) gives,
$\xi_{1}=\frac{|\mathscr{G}_{H}|^{2}}{|V(H)|^{2}}\left[\sum_{1\leq a,\,b\leq|V(H)|}t\left(H\bigoplus_{a,\,b}H,\,W\right)-|V(H)|^{2}t(H,\,W)^{2}\right].$
@@
Finally, invoking Bhattacharya et al. (2023, Lemma 5.4) we conclude that $\xi_{1}=0$ if and only if $W$ is $H$-regular. Thus, when $W$ is $H$-regular, $T(H,G_{n})$ is a first order degenerate $U$-statistics. Hence, by Lee (2019, Section 1.3, Theorem 3), $\mathrm{Var}[T(H,\,G_{n})]\lesssim_{H,\,W}\frac{1}{n^{2}}$. This completes the proof of Lemma 5.1. $\Box$
A.4. Proof of Proposition 6.1
Recall the definition of $T(H,G_{n})$ from (5.1). Since $W$ is not $H$-regular, by Lemma 5.1 $T(H,G_{n})$ is a $U$-statistic of order $|V(H)|$ which is non-degenerate. Hence, by Lee (2019, Section 1.3, Theorem 3), $\mathrm{Var}[T(H,G_{n})]\lesssim_{H,W}\frac{1}{n}$, and recalling (5.2),
Also, recalling (3.4) in the case when $W$ is not $H$-regular,
$\mathrm{Var}[\Delta(H,G_{n})]\asymp_{H,W}\max_{F\subseteq H:|V(F)|\geq 1}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}.$ (A.1)
We claim that when $n\rho_{n}^{m_{1}(H)}\ll 1,$ the maximum in (A.1) is obtained for some subgraph $F\subseteq H$ with $|V(F)|\geq 2.$ Indeed, when $n\rho_{n}^{m_{1}(H)}\ll 1,$ there exists a subgraph $F\subseteq H$ with $|V(F)|\geq 2$ such that $n^{|V(F)|-1}\rho_{n}^{|E(F)|}\ll 1$ or equivalently, $n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}\ll n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}.$ Thus,
$\frac{\mathrm{Var}[\Delta_{2}(H,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\lesssim_{H,W}\min_{F\subseteq H,|V(F)|\geq 2}n^{|V(F)|-1}\rho_{n}^{|E(F)|}\to 0,$
since $n\rho_{n}^{m_{1}(H)}\ll 1.$ This proves Proposition 6.1 (1).
Next, for the second part, from Lee (2019, Section 1.3, Theorem 3) we know that
$\lim_{n\to\infty}n\mathrm{Var}[T(H,G_{n})]=|V(H)|^{2}\xi_{1},$
$\mathrm{Var}[\Delta_{2}(H,G_{n})]=\frac{|V(H)|^{2}}{n}\binom{n}{|V(H)|}^{2}\rho_{n}^{2|E(H)|}\xi_{1}+o(n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}).$ (A.2)
Now, continuing from (3.6), we have
$\mathrm{Var}[\Delta(H,G_{n})]$ $=\sum_{F\subseteq H:|V(F)|=1}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq F\end{subarray}}\mathrm{Cov}[I_{s},I_{t}]+\sum_{F\subseteq H:|V(F)|\geq 2}\sum_{\stackrel{{\scriptstyle H_{s},H_{t}\in\mathcal{H}_{n}}}{{H_{s}\cap H_{t}\simeq F}}}\mathrm{Cov}[I_{s},I_{t}],$ (A.3)
where $F$ is the subgraph of $H$ isomorphic to $H_{s}\cap H_{t}=(V(H_{s})\cap V(H_{t}),E(H_{s})\cap E(H_{t}))$. Note that for a fixed subgraph $F\subseteq H,$ the number of pairs $(H_{s},H_{t})$ having $H_{s}\cap H_{t}$ isomorphic to $F$ is $O_{H}(n^{2|V(H)|-|V(F)|}).$ Hence, the second term in (A.3) can be bounded as,
$\sum_{F\subseteq H:|V(F)|\geq 2}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq F\end{subarray}}\mathrm{Cov}[I_{s},I_{t}]$ $\lesssim_{H,W}\sum_{F\subseteq H:|V(F)|\geq 2}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}$
$=o(n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}),$
since $n\rho_{n}^{m_{1}(H)}\gg 1$ is equivalent to the condition that for all subgraphs $F\subseteq H$ with $|V(F)|\geq 2,$ we have $n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}\ll n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}.$ To compute the first term in (A.3), first choose the vertex sets of the two copies of $H$ sharing a single vertex. This is equivalent to finding the number of ordered pairs of subsets $(A_{1},A_{2})$ of $[n]$ such that $|A_{1}\cap A_{2}|=1.$ Observe that the number of such ordered pairs is $|V(H)|\binom{n}{|V(H)|}\binom{n-|V(H)|}{|V(H)|-1}$. Once we have fixed the vertex sets of the two copies, we can trace the steps of the proof of Lemma 5.1 to obtain
$\mathrm{Var}[\Delta(H,G_{n})]=|V(H)|\binom{n}{|V(H)|}\binom{n-|V(H)|}{|V(H)|-1}\rho_{n}^{2|E(H)|}\xi_{1}+o(n^{2|V(H)|-1}\rho_{n}^{2|E(H)|}).$ (A.4)
from (A.3). Taking the ratio of (A.2) and (A.4) it follows that $\frac{\mathrm{Var}[\Delta_{2}(H,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\rightarrow 1$. Hence, recalling (2.3) the result in Proposition 6.1 (2) follows. $\Box$
A.5. Joint Convergence from Marginal Convergence